OpenStudy (anonymous):

A rectangle has a length of 9.00 inches and a width of 4.25 inches. What is the area in square inches? Round to the correct significant digits.

OpenStudy (texaschic101):

A = L * W A = 9 * 4.25 A = 38.25 I am not sure about rounding to significant digits... Not sure if it is : 38 or 38.3 or 38.25

OpenStudy (anonymous):

Alright, im going to try 38.3, I had gotten 38.2 but now that I start thinking about it 38.3 sounds more like it. Thank you!